please help solve
factor by grouping
4+xy-2y-2x
i am not sure how you know how to pair them up

Answer by Earlsdon(6294)

It really makes no difference which pair of terms you group together becuase you will get the same answer as long as you pair those terms that have the same variable, when possible:
Factor by grouping:
Example 1): Group the x-terms and the y-terms.
4+xy-2y-2x = (xy-2y)+(4-2x) Factor y from the first pair and 2 from the second pair:
y(x-2)+2(2-x) From the second term, factor -1 and rewrite it.
y(x-2)-2(x-2) Now factor the (x-2)
(x-2)(y-2) ...and that's it!
Let's try the other grouping:
4+xy-2y-2x = (xy-2x)+(4-2y) Factor x from the first pair and 2 from the second pair.
x(y-2)+2(2-y) From the second term, factor -1 and rewrite it.
x(y-2)-2(y-2) Now factor the (y-2)
(y-2)(x-2) or (x-2)(y-2)